OpenStudy (anonymous):

you draw 4 cards from a deck of cards, one after the other, without replacing them. What is the probability of drawing all 4 aces?

OpenStudy (anonymous):

How many aces are there in the deck of 52 cards ??

OpenStudy (anonymous):

There are 4 aces in the deck of 52 cards.. So, probability of selecting One Ace card will be : \[P(First \; Ace) = \frac{4}{52}\]

OpenStudy (anonymous):

Now that Ace has been taken and we have not replaced it in the deck: So now you are left with 3 Aces and 51 cards in total..

OpenStudy (anonymous):

Probability of choosing again an Ace will be: \[P(Second \; Ace) = \frac{3}{51}\] Similarly: \[P( Third \; Ace) = \frac{2}{50}\] And finally : \[P( Fourth \; Ace) = \frac{1}{49}\]

OpenStudy (anonymous):

So Total Probability of getting 4 aces will be : \[P = \frac{4}{52} \times \frac{3}{51} \times \frac{2}{50} \times \frac{1}{49} = Calculate\]

OpenStudy (anonymous):

Are you getting @sam66 ??

OpenStudy (anonymous):

i think so, so would the answer be 1/270725?

OpenStudy (anonymous):

I have not calculated it yet.. Wait...

OpenStudy (anonymous):

Yep I got the same...

OpenStudy (anonymous):

awesome, thanks sooo much

OpenStudy (anonymous):

Welcome dear...